# ALPy Cogenesis

### v1

@@ III Dirac ALPy cogenesis
$-\mathcal{L}$ $\supset\frac{\partial_{\mu}a}{f}\overline{N_{R_{i}}}(\lambda_{i\alpha}\gamma^{\mu}+\lambda^{\prime}_{i\alpha}\gamma^{\mu}\gamma^{5})\nu_{R_{\alpha}}$
@@
$\epsilon_{\nu_{R}}$ $=\frac{{\rm Im}[\widetilde{\lambda}^{*}_{j\alpha}\widetilde{h}_{Lj}\widetilde{h}^{*}_{Li}\widetilde{\lambda}_{i\alpha}]}{|\widetilde{h}_{Li}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\chi_{L}a)\frac{r_{ij}}{1-r^{2}_{ij}}$
$+\frac{{\rm Im}[\widetilde{\lambda}^{*}_{j\alpha}\widetilde{h}_{Rj}\widetilde{h}^{*}_{Ri}\widetilde{\lambda}_{i\alpha}]}{|\widetilde{h}_{Li}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\chi_{R}a)\frac{1}{1-r^{2}_{ij}},$ (21)
$\epsilon_{\chi_{R}}$ $=\frac{{\rm Im}[\widetilde{h}^{*}_{Rj}\widetilde{h}_{Lj}\widetilde{h}^{*}_{Li}\widetilde{h}_{Ri}]}{|\widetilde{h}_{Li}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\chi_{L}a)\frac{r_{ij}}{1-r^{2}_{ij}},$
$+\frac{{\rm Im}[\widetilde{h}^{*}_{Rj}\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}^{*}_{i\alpha}\widetilde{h}_{Ri}]}{|\widetilde{\lambda}_{i\alpha}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\nu_{R}a)\frac{1}{1-r^{2}_{ij}}$ (22)
$\epsilon_{\chi_{L}}$ $=\frac{{\rm Im}[\widetilde{h}^{*}_{Lj}\widetilde{h}_{Rj}\widetilde{h}^{*}_{Ri}\widetilde{h}_{Li}]}{|\widetilde{h}_{Ri}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\chi_{R}a)\frac{r_{ij}}{1-r^{2}_{ij}},$
$+\frac{{\rm Im}[\widetilde{h}^{*}_{Lj}\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}^{*}_{i\alpha}\widetilde{h}_{Li}]}{|\widetilde{\lambda}_{i\alpha}|^{2}\widetilde{\Gamma}_{\rm tot}}\Gamma_{0}(N_{i}\rightarrow\nu_{R}a)\frac{r_{ij}}{1-r^{2}_{ij}}$ (23)
@@
$\dfrac{d\eta_{\Delta\nu_{R}}}{dz}$ $=\frac{1}{s\mathcal{H}z}\bigg{[}\epsilon_{\nu_{R}}\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\nu_{R})$
@@
$\dfrac{d\eta_{\Delta\chi_{R}}}{dz}$ $=\frac{1}{s\mathcal{H}z}\bigg{[}\epsilon_{\chi_{R}}\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\chi_{R})$
@@
$\dfrac{d\eta_{\Delta\chi_{L}}}{dz}$ $=\frac{1}{s\mathcal{H}z}\bigg{[}\epsilon_{\chi_{L}}\left(\frac{\eta_{N_{1L}}}{\eta^{\rm eq}_{N_{1L}}}-1\right)\gamma(N_{1L}\rightarrow a\chi_{L})$
@@
$\dfrac{d\eta_{N_{1L}}}{dz}$ $=-\frac{1}{s\mathcal{H}z}\bigg{[}\left(\frac{\eta_{N_{1L}}}{\eta^{\rm eq}_{N_{1L}}}-1\right)\gamma(N_{1L}\rightarrow a\chi_{L})\bigg{]},$ (27)
$\dfrac{d\eta_{N_{1R}}}{dz}$ $=-\frac{1}{s\mathcal{H}z}\bigg{[}\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\chi_{R})$
$+\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\nu_{R})\bigg{]}.$ (28)
In the above equations, the thermally-averaged reaction densities for the $2-2$ scattering processes are defined as
[…]
where $\hat{\sigma}(s)$ is the reduced cross section for the process and is given by
@@
where $n_{i}^{\rm eq}$ and $n_{j}^{\rm eq}$ are the equilibrium number densities of species $i$ and $j$ respectively. For decay $A\longrightarrow B$ the reaction densities are defined as
@@
In order to prevent the asymmetry in $\nu_{R}$ from being washed out, one has to ensure that the processes $\nu_{R}a\leftrightarrow\chi a$ do not equilibrate. This can be taken care of by keeping the rates of these processes below the Hubble expansion rate of the Universe $\mathcal{H}$. This leads to the following constraint
@@
Figure 7: Allowed parameter space of Dirac ALPy cogenesis scenario in $f-M_{1}$ plane considering Dimensionless couplings $\sim 10^{-2}$. The shaded regions are disfavored due to different constraints, leaving only the white region allowed for Dirac ALPy cogenesis.
This leads to the generation of asymmetries in $\nu_{R}$ and $\chi$ as $N_{i}$ decays out-of-equilibrium. Similar to the minimal scenario discussed before, the asymmetry in $\nu_{R}$ gets transferred to left-handed lepton doublets via Yukawa couplings with the neutrinophilic Higgs doublet. The lepton asymmetry is then converted into baryon asymmetry via electroweak sphalerons. The asymmetry stored in $\chi$ can survive as asymmetric dark matter. Fig. 6 shows the evolution of comoving asymmetries in $\nu_{R},\chi_{R},\chi_{L}$ for fixed values of CP asymmetry parameters, heavy fermion mass and axion decay constant. For this particular benchmark, we choose $\epsilon_{\nu_{R}}=3\epsilon,\epsilon_{\chi_{R}}=-2\epsilon,\epsilon_{\chi_{L}}=-\epsilon$ with $\epsilon=2.256\times 10^{-7}$. While the net CP asymmetry is zero by unitarity, the asymmetry generated in $\nu_{R}$ is consistent with the required lepton asymmetry to generate the observed BAU. Fig. 7 shows the parameter space of Dirac ALPy cogenesis in $f-M_{1}$ plane. The solid contours correspond to leptogenesis allowed parameter space for two different values of CP asymmetry parameter. Similar to the minimal scenario, the shaded region labeled as $\Gamma_{\rm agg}/\mathcal{H}<1$ corresponds to axions not being part of the bath. The blue shaded region is ruled out due to strong washouts in equilibrium. The horizontal shaded region at the top leads to axion DM overproduction via vacuum misalignment.
Depending upon axion parameter space, $\chi$ can constitute the observed DM entirely or partially. The net dark asymmetry adds up to generate an asymmetric $\chi$ abundance at present epoch $(T=T_{0})$ given by
@@
where $S$ is the same dilution factor used in Eq. (17) and $\rho_{\rm cr}(T_{0})=\frac{3\mathcal{H}^{2}_{0}}{8\pi G}$ is the critical density of the Universe at present. Since $\lvert\eta_{\Delta\chi_{L}}+\eta_{\Delta\chi_{R}}\rvert\sim\eta_{\nu_{R}}$ and $\Omega_{\rm DM}\approx 5.36\Omega_{B}$ Planck:2018vyg , we get $m_{\chi}\leq 1.9m_{p}$ such that $\Omega_{\chi}\leq\Omega_{\rm DM}$. However, $\chi$ is not perfectly stable and can have three-body decay with decay width given by
@@
The green shaded region at the bottom left corner of Fig. 7 shows the region of parameter space in $f-m_{\chi}$ plane where $\chi$ lifetime is less than the age of the Universe. Clearly, for realistic values of $f$, particularly for QCD axion, lifetime criteria for $\chi$ is trivially satisfied. On the other hand, the red shaded region of Fig. 7 corresponds to overproduced relic of $\chi$. In order to constrain the parameter space in a conservative way, here we consider $m_{\chi}$ to be same as the one-loop mass given by
@@
This leaves only the white region of Fig. 7 allowed from successful Dirac ALPy cogenesis and dark matter abundance. It should be noted that the model also allows one-loop $\chi_{L}-\nu_{R}$ mixing $m^{\rm loop}_{\chi\nu}$. However, in the limit of sub-eV Dirac neutrino mass $M_{D}\ll m_{\chi},m^{\rm loop}_{\chi\nu}$, it does not lead to additional constraints from neutrino data or $\chi$ lifetime requirements.
Figure 8: Axion-photon coupling versus axion mass satisfying correct baryon asymmetry for minimal (left) and Dirac (right) ALPy cogenesis. The colored circular points correspond to leptogenesis allowed parameter space for $\epsilon_{1}=3.67\times 10^{-7}$ (minimal) and $\epsilon=2.26\times 10^{-7}$ (Dirac) with color bar indicating the scale of leptogenesis $M_{1}$. The shaded regions correspond to different bounds or sensitivities. The cyan colored $\star$ corresponds to the point where QCD axion DM satisfies observed relic while the cyan colored solid line (right panel) corresponds to the region where axion and $\chi$ together generate the observed DM relic.
@@ IV Detection Aspects
As discussed above, leptogenesis in our setup depends upon the heavy fermion mass $M_{1}$ as well as the axion decay constant $f$. This opens up the possibility to probe ALPy cogenesis favored parameter space indirectly at axion detection experiments. Fig. 8 shows the parameter space in the plane of axion-photon coupling $g_{a\gamma}$ and axion mass $m_{a}$. The current experimental bounds on the axion-photon coupling from various experiments or observables are shown by the solid color lines (CAST CAST:2007jps ; CAST:2017uph , SN87A PhysRevLett.60.1797 ; PhysRevD.39.1020 ; PhysRevLett.60.1793 , NGC 1275 Fermi-LAT:2016nkz , ADMX ADMX:2006kgb ; Stern:2016bbw ; ADMX:2019uok , Globular clusters Ayala:2014pea ) whereas future experimental sensitivities or observables are shown by the dashed lines (CASPEr Budker:2013hfa , FLASH Alesini:2017ifp ; Alesini:2019nzq ; Alesini:2023qed , ABRACADABRA Kahn:2016aff ; Ouellet:2018beu , CULTASK Lee:2020cfj ; Semertzidis:2019gkj , MADMAX Caldwell:2016dcw , IAXO Vogel:2013bta ; IAXO:2019mpb , Fermi-LAT Meyer:2016wrm , BH superradiance Cardoso:2018tly ). While for ALP, axion-photon coupling is not related to axion mass, it is not so for QCD axion. The solid red line in Fig. 8 shows $g_{a\gamma}-m_{a}$ relation for QCD axion considering the PQ model to be of Kim-Shifman-Vainshtein-Zakharov (KSVZ) Kim:1979if ; Shifman:1979if type. The circular points on this line correspond to leptogenesis allowed parameter space with the color bar indicating the scale $M_{1}$. Assuming axion DM to be produced from vacuum misalignment, the minimal ALPy cogenesis is consistent with the observed BAU and DM for the point denoted as cyan colored $\star$ in the left panel plot of Fig. 8. The parameter space widens up to the cyan colored line in the right panel plot of Fig. 8 in Dirac ALPy cogenesis as deficit in axion DM abundance can be filled by asymmetric dark fermion $\chi$. This keeps the Dirac ALPy cogenesis verifiable across a wider range of experiments compared to the minimal scenario.
@@
There exist other detection prospects of our scenario in addition to the ones outlined in Fig. 8. For the choice of our parameters, axions are also produced thermally in the early Universe leaving a hot axion component in addition to the cold axion DM part. For sub-eV axion mass, this leads to component of dark radiation which can be probed via future CMB measurements of effective relativistic degrees of freedom $N_{\rm eff}$. The current bound on such excess dark radiation or $\Delta N_{\rm eff}$ from Planck 2018 measurement reads $\Delta N_{\rm eff}<0.285$ at $2\sigma$ Planck:2018vyg . Future CMB experiments like CMB-S4 and CMB-HD are sensitive to $\Delta N_{\rm eff}=0.06$ Abazajian:2019eic and $\Delta N_{\rm eff}=0.014$ CMB-HD:2022bsz respectively. A thermalised particle like axion with one degree of freedom which decouples from the bath above electroweak scale temperature leads to $\Delta N_{\rm eff}=0.027$ Borah:2024kfn . In the Dirac ALPy cogenesis scenario, we also have additional contribution to dark radiation from light Dirac neutrinos. Thermalised light Dirac neutrinos with decoupling temperature above electroweak scale leads to $\Delta N_{\rm eff}=0.14$ Abazajian:2019oqj . Therefore, minimal ALPy cogenesis can be probed at future CMB experiments like CMB-HD CMB-HD:2022bsz while Dirac ALPy cogenesis remains within reach of multiple CMB missions due to enhanced $\Delta N_{\rm eff}$.
@@
While we have shown the detection prospects of our leptogenesis scenario at axion search and CMB experiments, there are other possible avenues as well. We briefly comment on them without going into the details. Firstly, the minimal ALPy cogenesis scenario has additional heavy neutral leptons (HNL) in sub-TeV to sub-GeV scales due to the Majorana mass of $\nu_{R}$. Since they mix with active neutrinos, they can have interesting experimental signatures, a recent review of which can be found in Abdullahi:2022jlv . HNL below electroweak scale can lead to a variety of final states depending upon their masses and mixing with active neutrinos. There are dedicated ongoing experiments like NA62 NA62:2017qcd , FASER FASER:2018eoc and upcoming experiments like DUNE DUNE:2020fgq , Codex-b CODEX-b:2019jve , MATHUSLACurtin:2018mvb to observe such signatures of such light HNL. Additionally, a keV scale HNL can also be a good DM candidate Drewes:2016upu . If the scale of such sterile neutrino is even lower, they can also have other interesting implications in neutrino oscillations, neutrinoless double beta decay, astrophysics etc. as summarised in deGouvea:2006gz . Secondly, thermal axions at the eV scale also constitute hot dark matter having other indirect detection prospects, as pointed out recently in Dror:2024ibf . On the other hand, the long-lived nature of asymmetric DM $\chi$ in Dirac ALPy cogenesis setup can also be probed at indirect search experiments looking for excess photons. Such excess photons can arise via $\chi\rightarrow aa\nu_{R}$ followed by axion decay or conversion into photons. Finally, both the models discussed here rely on an additional Higgs doublet which can have a rich phenomenology Branco:2011iw .

### v2

@@ III Dirac ALPy cogenesis
-$\mathcal{L}$ $\supset\frac{\partial_{\mu}a}{f}\overline{N_{R_{i}}}(\lambda_{i\alpha}\gamma^{\mu}+\lambda^{\prime}_{i\alpha}\gamma^{\mu}\gamma^{5})\nu_{R_{\alpha}}$
@@
$\epsilon_{\nu_{R}}$ $=\frac{{\rm Im}[\widetilde{\lambda}^{*}_{j\alpha}\widetilde{h}_{Lj}\widetilde{h}^{*}_{Li}\widetilde{\lambda}_{i\alpha}]}{|\widetilde{h}_{Li}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\chi_{L}a)\frac{r_{ij}}{1-r^{2}_{ij}}$
$+\frac{{\rm Im}[\widetilde{\lambda}^{*}_{j\alpha}\widetilde{h}_{Rj}\widetilde{h}^{*}_{Ri}\widetilde{\lambda}_{i\alpha}]}{|\widetilde{h}_{Ri}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\chi_{R}a)\frac{1}{1-r^{2}_{ij}},$ (23)
$\epsilon_{\chi_{R}}$ $=\frac{{\rm Im}[\widetilde{h}^{*}_{Rj}\widetilde{h}_{Lj}\widetilde{h}^{*}_{Li}\widetilde{h}_{Ri}]}{|\widetilde{h}_{Li}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\chi_{L}a)\frac{r_{ij}}{1-r^{2}_{ij}}$
$+\frac{{\rm Im}[\widetilde{h}^{*}_{Rj}\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}^{*}_{i\alpha}\widetilde{h}_{Ri}]}{|\widetilde{\lambda}_{i\alpha}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\nu_{R}a)\frac{1}{1-r^{2}_{ij}},$ (24)
$\epsilon_{\chi_{L}}$ $=\frac{{\rm Im}[\widetilde{h}^{*}_{Lj}\widetilde{h}_{Rj}\widetilde{h}^{*}_{Ri}\widetilde{h}_{Li}]}{|\widetilde{h}_{Ri}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\chi_{R}a)\frac{r_{ij}}{1-r^{2}_{ij}}$
$+\frac{{\rm Im}[\widetilde{h}^{*}_{Lj}\widetilde{\lambda}_{j\alpha}\widetilde{\lambda}^{*}_{i\alpha}\widetilde{h}_{Li}]}{|\widetilde{\lambda}_{i\alpha}|^{2}\widetilde{\Gamma}_{\rm tot}M_{i}}\Gamma_{0}(N_{i}\rightarrow\nu_{R}a)\frac{r_{ij}}{1-r^{2}_{ij}},$ (25)
@@
$\dfrac{d\eta_{\Delta\nu_{R}}}{dz}$ $=\frac{1}{n_{\gamma}\mathcal{H}z}\bigg[\epsilon_{\nu_{R}}\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\nu_{R})$
@@
$\dfrac{d\eta_{\Delta\chi_{R}}}{dz}$ $=\frac{1}{n_{\gamma}\mathcal{H}z}\bigg[\epsilon_{\chi_{R}}\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\chi_{R})$
@@
$\dfrac{d\eta_{\Delta\chi_{L}}}{dz}$ $=\frac{1}{n_{\gamma}\mathcal{H}z}\bigg[\epsilon_{\chi_{L}}\left(\frac{\eta_{N_{1L}}}{\eta^{\rm eq}_{N_{1L}}}-1\right)\gamma(N_{1L}\rightarrow a\chi_{L})$
@@
$\dfrac{d\eta_{N_{1L}}}{dz}$ $=-\frac{1}{n_{\gamma}\mathcal{H}z}\bigg[\left(\frac{\eta_{N_{1L}}}{\eta^{\rm eq}_{N_{1L}}}-1\right)\gamma(N_{1L}\rightarrow a\chi_{L})\bigg],$ (29)
$\dfrac{d\eta_{N_{1R}}}{dz}$ $=-\frac{1}{n_{\gamma}\mathcal{H}z}\bigg[\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\chi_{R})$
$+\left(\frac{\eta_{N_{1R}}}{\eta^{\rm eq}_{N_{1R}}}-1\right)\gamma(N_{1R}\rightarrow a\nu_{R})\bigg],$ (30)
where $\eta_{x}=n_{x}/n_{\gamma}$, $\eta_{\Delta x}=(n_{x}-n_{\bar{x}})_{/}n_{\gamma}$ and $z=M_{1}/T$. In the above equations, the thermally-averaged reaction densities for the 222-2 scattering processes are defined as
$\gamma_{ij\longrightarrow kl}=\frac{T}{64\pi^{4}}\int_{(m_{i}+m_{j})^{2}}^{\infty}ds\sqrt{s}K_{1}\left(\sqrt{s}/T\right)\hat{\sigma}(s),$ (31)
where $\hat{\sigma}(s)$ is the reduced cross-section for the process and is given by
@@
where $n_{i}^{\rm eq}$ and $n_{j}^{\rm eq}$ are the equilibrium number densities of species $i$ and $j$ respectively. For the decay $A\longrightarrow B$, the reaction density is defined as
@@
In order to prevent the asymmetry in $\nu_{R}$ from being washed out, one has to ensure that the process $\nu_{R}a\leftrightarrow\chi a$ does not equilibrate. This can be taken care of by keeping the rates of this process below the Hubble expansion rate of the Universe $\mathcal{H}$. This leads to the following constraint
@@
Figure 7: Allowed parameter space of Dirac ALPy cogenesis scenario in $f-M_{1}$ plane considering Dimensionless couplings $\tilde{\lambda}_{i\alpha},\tilde{h}_{Li},\tilde{h}_{Ri}\sim 10^{-2}$. The shaded regions are disfavored due to different constraints, leaving only the white region allowed for Dirac ALPy cogenesis.
This leads to the generation of asymmetries in $\nu_{R}$ and $\chi$ as $N_{i}$ decays out-of-equilibrium. Similar to the minimal scenario discussed before, the asymmetry in $\nu_{R}$ gets transferred to the left-handed lepton doublets via Yukawa couplings with the neutrinophilic Higgs doublet. The lepton asymmetry is then converted into baryon asymmetry via the electroweak sphalerons. The asymmetry stored in $\chi$ can survive as asymmetric dark matter. Fig. 6 shows the evolution of comoving asymmetries in $\nu_{R},\chi_{R},\chi_{L}$ for fixed values of the CP asymmetry parameters, heavy fermion mass and axion decay constant. For this particular benchmark, we choose $\epsilon_{\nu_{R}}=3\epsilon,\epsilon_{\chi_{R}}=-2\epsilon,\epsilon_{\chi_{L}}=-\epsilon$ with $\epsilon=2.256\times 10^{-7}$. While the net CP asymmetry is zero by unitarity, the asymmetry generated in $\nu_{R}$ is consistent with the required lepton asymmetry to generate the observed BAU. Fig. 7 shows the parameter space of Dirac ALPy cogenesis in $f-M_{1}$ plane. The solid contours correspond to the leptogenesis allowed parameter space for two different values of the CP asymmetry parameter. The points represented by the solid blue and maroon colored contours correspond to the parameter space consistent with the observed baryon asymmetry for CP asymmetry parameter values of $\epsilon=2.26\times 10^{-7},\epsilon=10^{-2}$ respectively. Similar to the minimal scenario, the shaded region labeled as $\Gamma_{\rm agg}/\mathcal{H}<1$ corresponds to axions not being part of the bath. In this region, calculation of baryon asymmetry requires solving the Boltzmann equation for axions and estimating the non-thermal abundance of heavy neutral fermions, which we do not pursue here. The blue shaded region is ruled out due to strong washouts in equilibrium. The horizontal shaded region at the top leads to axion DM overproduction via vacuum misalignment. Similar to Fig. 4, here also the dimensionless couplings $\tilde{\lambda}_{i\alpha},\tilde{h}_{Li},\tilde{h}_{Ri}$ are assumed to be $\sim 10^{-2}$. The constraints from washouts, $\chi$ relic and $\chi$ lifetime become stronger for larger values of these dimensionless couplings, shrinking the allowed white colored region in Fig. 7.
Depending upon the axion parameter space, $\chi$ can constitute the observed DM entirely or partially. The net dark asymmetry adds up to generate an asymmetric $\chi$ abundance at present epoch $(T=T_{0})$ given by
@@
where $S$ is the same dilution factor used in Eq. (18) and $\rho_{\rm cr}(T_{0})=\frac{3\mathcal{H}^{2}_{0}}{8\pi G}$ is the critical density of the Universe at present. Since $\lvert\eta_{\Delta\chi_{L}}+\eta_{\Delta\chi_{R}}\rvert\sim\eta_{\Delta\nu_{R}}$ and $\Omega_{\rm DM}\approx 5.36\Omega_{B}$ Planck:2018vyg , we get $m_{\chi}\leq 1.9m_{p}$ such that $\Omega_{\chi}\leq\Omega_{\rm DM}$. However, $\chi$ is not perfectly stable and can have three-body decay with decay width given by
@@
In order to constrain the parameter space in a conservative way, here we consider $m_{\chi}$ to be same as the one-loop mass given by
@@
The green shaded region of Fig. 7 shows the region of parameter space in $f-m_{\chi}$ plane where $\chi$ lifetime $\tau_{\chi}$ is less than the age of the Universe. Clearly, for realistic values of $f$, particularly for the QCD axion, lifetime criteria for $\chi$ rules out the high mass regime of $N_{1}$, shown by the green shaded part. On the other hand, the red shaded region of Fig. 7 corresponds to the overproduced relic of $\chi$. This leaves only the white region of Fig. 7 allowed from successful Dirac ALPy cogenesis and dark matter abundance. It should be noted that the model also allows one-loop $\chi_{L}-\nu_{R}$ mixing $m^{\rm loop}_{\chi\nu}$. However, in the limit of sub-eV Dirac neutrino mass $M_{D}=\frac{yv_{2}}{\sqrt{2}}\ll m_{\chi},m^{\rm loop}_{\chi\nu}$, it does not lead to additional constraints from neutrino data or $\chi$ lifetime requirements.
Figure 8: Axion-photon coupling versus axion mass satisfying correct baryon asymmetry for minimal (left) and Dirac (right) ALPy cogenesis. The colored circular points correspond to leptogenesis allowed parameter space for $\epsilon_{1}=3.67\times 10^{-7}$ (minimal) and $\epsilon=2.26\times 10^{-7}$ (Dirac) with color bar indicating the scale of leptogenesis $M_{1}$. The other parameters are kept same as in Fig. 4 and Fig. 7. The shaded regions correspond to different bounds or sensitivities. The cyan colored $\star$ corresponds to the point where QCD axion DM produced via vacuum misalignment satisfies observed relic while the cyan colored solid line (right panel) corresponds to the region where axion and $\chi$ together generate the observed DM relic. The pink and gray colored bands on the right panel correspond to short-lived $\chi$ compared to the lifetime of the Universe $\tau_{U}$ and overproduced $\chi$, respectively.
@@ IV Detection Aspects
As discussed above, leptogenesis in our setup depends upon the heavy fermion mass $M_{1}$ as well as the axion decay constant $f$. This opens up the possibility to probe ALPy cogenesis favored parameter space indirectly at axion detection experiments. Fig. 8 shows the parameter space in the plane of axion-photon coupling $g_{a\gamma}$ and axion mass $m_{a}$. The current experimental bounds on the axion-photon coupling from various experiments or observables are shown by the solid color lines (CAST CAST:2007jps ; CAST:2017uph , SN87A PhysRevLett.60.1797 ; PhysRevD.39.1020 ; PhysRevLett.60.1793 , NGC 1275 Fermi-LAT:2016nkz , ADMX ADMX:2006kgb ; Stern:2016bbw ; ADMX:2019uok , Globular clusters Ayala:2014pea ) whereas future experimental sensitivities or observables are shown by the dashed lines (CASPEr Budker:2013hfa , FLASH Alesini:2017ifp ; Alesini:2019nzq ; Alesini:2023qed , ABRACADABRA Kahn:2016aff ; Ouellet:2018beu , CULTASK Lee:2020cfj ; Semertzidis:2019gkj , MADMAX Caldwell:2016dcw , IAXO Vogel:2013bta ; IAXO:2019mpb , Fermi-LAT Meyer:2016wrm , BH superradiance Cardoso:2018tly ). While for ALP, the axion-photon coupling is not related to axion mass, it is not so for QCD axion. The solid red line in Fig. 8 shows the $g_{a\gamma}-m_{a}$ relation for QCD axion considering the PQ model to be of Kim-Shifman-Vainshtein-Zakharov (KSVZ) Kim:1979if ; Shifman:1979if type. The circular points on this line correspond to leptogenesis allowed parameter space with the color bar indicating the scale $M_{1}$. Assuming axion DM to be produced from vacuum misalignment, the minimal ALPy cogenesis is consistent with the observed BAU and DM for the point denoted as cyan colored $\star$ in the left panel plot of Fig. 8. The parameter space widens up to the cyan colored line in the right panel plot of Fig. 8 in Dirac ALPy cogenesis as the deficit in axion DM abundance can be filled by asymmetric dark fermion $\chi$. This keeps the parameter space consistent with Dirac ALPy cogenesis verifiable across a wider range of experiments compared to the minimal scenario. In both the plots shown in Fig. 8, we have chosen the values of the CP asymmetry parameter in such a way that a wide range of $f$ remains allowed leading to a wider range of $g_{a\gamma}-m_{a}$ parameter space within reach of different experiments. For larger values of the CP asymmetry parameter say, $10^{-2}$, $f$ is restricted to small values of $f\lesssim 10^{6}$ GeV (as seen from Fig. 4 and Fig. 7) which are already ruled out for the QCD axions. For intermediate values of the CP asymmetry parameter, the allowed colored band of circular points in Fig. 8 will get squeezed to higher values of $g_{a\gamma},m_{a}$.
@@
There exist other detection prospects of our scenario in addition to the ones outlined in Fig. 8. For the choice of our parameters, axions are also produced thermally in the early Universe leaving a hot axion component in addition to the cold axion DM part. For sub-eV axion mass, this leads to component of dark radiation which can be probed via future CMB measurements of effective relativistic degrees of freedom $N_{\rm eff}$. The current bound on such excess dark radiation or $\Delta N_{\rm eff}$ from Planck 2018 measurement reads $\Delta N_{\rm eff}<0.285$ at $2\sigma$ Planck:2018vyg . Future CMB experiments like CMB-S4 and CMB-HD are sensitive to $\Delta N_{\rm eff}=0.06$ Abazajian:2019eic and $\Delta N_{\rm eff}=0.014$ CMB-HD:2022bsz respectively. A thermalized particle like axion with one degree of freedom which decouples from the bath above electroweak scale temperature leads to $\Delta N_{\rm eff}=0.027$ Borah:2024kfn . In the Dirac ALPy cogenesis scenario, we also have additional contribution to dark radiation from light Dirac neutrinos. Thermalized light Dirac neutrinos with decoupling temperature above electroweak scale leads to $\Delta N_{\rm eff}=0.14$ Abazajian:2019oqj . Therefore, minimal ALPy cogenesis can be probed at future CMB experiments like CMB-HD CMB-HD:2022bsz while Dirac ALPy cogenesis remains within reach of multiple CMB missions due to enhanced $\Delta N_{\rm eff}$.
@@
While we have shown the detection prospects of our leptogenesis scenario at axion search and CMB experiments, there are other possible avenues as well. We briefly comment on them without going into the details. Firstly, the minimal ALPy cogenesis scenario has additional heavy neutral leptons (HNL) in sub-TeV to sub-GeV scales due to the Majorana mass of $\nu_{R}$. Since they mix with active neutrinos, they can have interesting experimental signatures, a recent review of which can be found in Abdullahi:2022jlv . HNL below electroweak scale can lead to a variety of final states depending upon their masses and mixing with active neutrinos. There are dedicated ongoing experiments like NA62 NA62:2017qcd , FASER FASER:2018eoc and upcoming experiments like DUNE DUNE:2020fgq , Codex-b CODEX-b:2019jve , MATHUSLACurtin:2018mvb to observe such signatures of such light HNL. Additionally, a keV scale HNL can also be a good DM candidate Drewes:2016upu . If the scale of such sterile neutrino is even lower, they can also have other interesting implications in neutrino oscillations, neutrinoless double beta decay, astrophysics etc. as summarized in deGouvea:2006gz . Secondly, thermal axions at the eV scale also constitute hot dark matter having other indirect detection prospects, as pointed out recently in Dror:2024ibf . On the other hand, the long-lived nature of asymmetric DM $\chi$ in Dirac ALPy cogenesis setup can also be probed at indirect search experiments looking for excess photons. Such excess photons can arise via $\chi\rightarrow aa\nu_{R}$ followed by axion decay or conversion into photons. Finally, both the models discussed here rely on an additional neutrinophilic Higgs doublet which can have a rich phenomenology Maitra:2014qea ; Haba:2011nb ; Gabriel:2008es ; Seto:2015rma ; Bertuzzo:2015ada ; Huitu:2017vye .